OpenStudy (anonymous):

PLEASE HELP AND SHOW ALL WORK: List the first four terms of a geometric sequence with T1 = 4 and Tn = -3Tn-1.

OpenStudy (anonymous):

t1= 4 \[Tn = (-3)^{n-1}\] Like that ?

OpenStudy (anonymous):

no like Tn= -3t^(n-1)

OpenStudy (primeralph):



OpenStudy (anonymous):

wait i got it wrong like this

OpenStudy (campbell_st):

well its saying to get the next term multiply the previous turn by -3 so if the 1st term is 4 then 2nd term is T2 = -3 x 4 a negative value 3rd term T3 = -3 x T3 a positive value 4th term T4 = -4 x T3 a negative value

OpenStudy (campbell_st):

oops T3 = -3 x T2

OpenStudy (campbell_st):

and T4 = -3 x T3

OpenStudy (campbell_st):

sorry for the typos

OpenStudy (anonymous):

But i'm confused over what Tn=-3Tn-1 means like how to substitute it or something and its fine

OpenStudy (campbell_st):

ok... what they needed to say is that n >= 2 so you know \[T_{1} = 4\] now let n = 2 using the formula \[T_{2} = -3 \times T_{2 -1}...so.... T_{2} = -3 \times T_{1}\] and when n = 3 \[T_{3} = -3 \times T_{3 -1}\] which is \[T_{3} = -3 \times T_{2}\] so all the subscript is doing is telling you the position of the term in the sequence

OpenStudy (anonymous):

oh that makes more sense so you're basically just multiplying by -3 to find each term?

OpenStudy (campbell_st):

thats correct...

OpenStudy (anonymous):

ok thank you so much

OpenStudy (campbell_st):

glad to help